OpenStudy (anonymous):

How do I factorise a polynomial like: v^4-16?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge a^2-b^2 = (a+b)(a-b)$$

OpenStudy (skullpatrol):

$$\Huge 16=4^2$$

OpenStudy (skullpatrol):

$$\Huge v^4=(v^2)^2$$

OpenStudy (skullpatrol):